OpenStudy (bloodydiamond):

derivative of y=(x-1)^2+1

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = \frac{ d }{ dx }(x-1)^2 + \frac{ d }{ dx }(1) = 2(x-1)(\frac{ d}{ dx }(x-1)) + 0 = 2(x-1)\]

OpenStudy (dariusx):

y = x^2 - 2x + 1 +1 y = x^2 -2x + 2 y' = 2x - 2